OpenStudy (anonymous):

Consider two polynomials P(x)=4x^(3)+ax^(2)+bx-5 and Q(x)=x^(2)+1. It is given that P(x) is divisible by x+1 and 2x+1. (a) Find the values of a and b. (b) Find the quotient and the remainder when P(x) is divided by Q(x). (c) Solve (8x-1)Q(x)-2P(x)-9=0.

OpenStudy (anonymous):

The divisibility by x+1 and 2x+1 gives you two roots. So you know P(roots)=0 so now you have a system of equations. Two equations two unknows. You can solve for a and b

OpenStudy (anonymous):

4(-1)^(3)+a(-1)^(2)+b(-1)-5 = 0 4(-1/2)^(3)+a(-1/2)^(2)+b(-1/2)-5 = 0

OpenStudy (anonymous):

4(-1)^(3)+a(-1)^(2)+b(-1)-5 = 0 4(-1/2)^(3)+a(-1/2)^(2)+b(-1/2)-5 = 0 a-b=9---(1) a-2b=-22---(2) i can't get the answer

OpenStudy (anonymous):

That 22 should be positive.

OpenStudy (anonymous):

ok got it.